OpenStudy (anonymous):

What is the slope for x=2- 3/2y?

OpenStudy (anonymous):

-2/3

OpenStudy (anonymous):

How do you get that answer?

OpenStudy (anonymous):

x = (2-3) / 2y is this the problem?

OpenStudy (anonymous):

x=2- 3/2 y y=-2/3 x+ 4/3

OpenStudy (anonymous):

y=mx+c where m is the slope.. compare these

OpenStudy (anonymous):

No it is x= 2+ (-3/2y)

OpenStudy (anonymous):

then it is not a linear equation i think ..

OpenStudy (anonymous):

would it be y= -2/3x+2? and the slope is 2/3?

OpenStudy (phi):

You have x=2- (3/2)y The easiest way to find the slope is put the equation in the form y = m x + b so add +(3/2)y to both sides (3/2)y + x = 2 add -x to both sides (3/2) y= -x + 2 multiply both sides by 2/3 y= -(2/3) x + 4/3 now match up with y= m x + b we see m= -2/3 . That is the slope. And, btw, the y-intercept is 4/3

OpenStudy (anonymous):

Thank you! So then if you have 2 slopes one -3/2 and the other -2/3 is that a perpendicular line?

OpenStudy (anonymous):

lol i thought the problem is x= 2+ (-3/2y) but its x = 2 - 3y/2 ..

OpenStudy (anonymous):

thats why i said earlier its not a linear ^^

OpenStudy (phi):

perpendicular lines have slopes m and -1/m so if you have -3/2 the other would have to be +2/3

OpenStudy (anonymous):

No, its x=2+(-3/2y).

OpenStudy (anonymous):

So then the line would be neither parellel or perpendicular?

OpenStudy (phi):

If you have another line and you correctly found its slope, then yes, neither.

OpenStudy (anonymous):

@lcombs the problem should be like this x = 2 + (-3/2)y

OpenStudy (anonymous):

Yes, it was a two slope problem with the first slope being -3/2 and the one we just worked out being -2/3.

OpenStudy (phi):

Then neither. Negative slope means the line is going down and to the right